A body of mass 8 kg is moving at 2 m/s without any external force on it. An internal explosion breaks the body into two equal parts and 16 joules of translational kinetic energy is imparted to the system by the explosion. Determine the speed and direction of motion of the blocks after the explosion

Grade:11

1 Answers

jitender
6 years ago
When a body of mass m travels in straight line motion with velocity u then if there is no external force acting on the body....And if explosion raje place due to internal forces then the momentum before and after explosion is reserved . And If body is divided into two equal parts then both the parts move in opposite direction in a direction perpendicular to the initial motion .Mu = mv + mvm= M/28×2 = 2×4×vv=2m/s